The ordered pair (-2,-1) is a solution of which of the following?
a. 4x+y=7
b. -4x-=7
c. 4x+y=-7
d. 4x-y=-7
e. -4x+y=-7

Answer by richwmiller(17219)

Just plug in (-2,-1) each of the equations and see which one comes up with an answer that is equal
(-2,-1) means x=-2 and y=-1
for example
2x-6y=5
2(-2)-6(-1)=5
-4+7=5
3=5
3 is not equal to 5 so(-2,-10 is not on the line and is not a solution to 2x-6y=5
The ordered pair that works for you will end up either 7=7 or -7=-7